You are evaluating the purchase for a manufacturing machine; the price is $120,000 and it will save company $50,000 of expenses every year. The machine is depreciated using three-year straight-line schedule and will be sold after that for $50,000. The project would have no effect on revenues, and the corporate tax rate is 30%. What is the project’s after-tax cash flow in Year 3?

(A)

$35,000.

(B)

$47,000.

(C)

$60,000.

(D)

$70,000.

(E)

$82,000.

2. Using the same information from previous question (Question 15). The machine will be sold for $50,000 after three years. What is the Year 3 after-tax cash flow if the machine is depreciated using MACRS 3-year class where the applicable depreciation rates are 33%, 45%, 15% and 7%.

(A)

$26,600.

(B)

$61,600.

(C)

$69,520.

(D)

$75,400.

(E)

$77,920.
Answer #1
1) (E) $       82,000
Working:
Step-1:Calculation of after tax operating cash flow
Saving of expense $       50,000
Depreciation $     -40,000
Profit before tax $       10,000
Tax Expense $        -3,000
Net Income $         7,000
Depreciation $       40,000
After tax operating cash flow $       47,000
Step-2:After tax cash flow from sale of asset
After tax cash flow from sale of asset = Before tax sale *(1- Tax rate)
= 50000*(1-0.30)
= $       35,000
Working:
Depreciation expense = Cost / Useful Life
= 120000/3
= $       40,000
Step-3:After tax cash flow in year 3
Operating cash flow $       47,000
Terminal Cash flow $       35,000
Cash flow in year 3 $       82,000
2) (E) $       77,920
Working:
Step-1:Calculation of after tax operating cash flow
Saving of expense $       50,000
Depreciation $     -18,000
Profit before tax $       32,000
Tax Expense $        -9,600
Net Income $       22,400
Depreciation $       18,000
After tax operating cash flow $       40,400
Step-2:After tax cash flow from sale of asset
After tax cash flow from sale of asset = Before tax sale *(1- Tax rate)
= 50000-((50000-8400)*0.30)
= $       37,520
Working:
Depreciation Schedule:
Year Cost Depreciation rate Depreciation expense Accumulated Depreciation expense Book Value
1 $       1,20,000 33% $ 39,600 $     39,600 $ 80,400
2 $       1,20,000 45% $ 54,000 $     93,600 $ 26,400
3 $       1,20,000 15% $ 18,000 $ 1,11,600 $    8,400
Step-3:After tax cash flow in year 3
Operating cash flow $       40,400
Terminal Cash flow $       37,520
Cash flow in year 3 $       77,920